# Matryoshka: Learning to Drive Black-Box LLMs with LLMs

### v1

@@ 3.2 Data Collection by Interacting with Black-Box LLM Environment
In the MDP formulation, we consider the action space as the set of possible guidance that can enhance the capabilities of black-box LLMs. The observation space is determined by the oracle evaluation function for each task, defined as $\text{eval}(\cdot)$, where the sampled supervision signal is denoted as $z$, with $z=1$ indicating that $f_{\theta}(x)$ is positive guidance while $z=0$ indicating $f_{\theta}(x)$ negative guidance. During the multi-turn interactions, if the observation $o_{t}$ at the $t$-th step returns a negative signal, the next action step $a_{t+1}$ involves modifying the intermediate guidance based on the feedback. The interactions continue until a positive signal is observed or the maximum number of interaction turns $T$ is reached.
For each input $x_{i}$, we perform $T$-step multi-turn interactions with the black-box LLM-based environment to obtain the trajectories $(a_{i,1},o_{i,1},a_{i,2},o_{i,2},\cdots,a_{i,T},o_{i,T})$. To increase the diversity of intermediate generations, we introduce randomness into the policy and repeat the entire interaction process $K$ times. This results in $K$ trajectories, yielding intermediate generations $\{g_{i,1},g_{i,2},\cdots,g_{i,K\times T}\}$ along with their corresponding observations $\{o_{i,1},o_{i,2},\cdots,o_{i,K\times T}\}$, which serve as sampling signals. We then sample the positive guidance $g_{i}^{+}$ from the set of guidance with positive observations, $g_{i}^{+}\sim\{g_{i,j}|o_{i,j}=1\}$ and the negative guidance from the remaining generations, $g_{i}^{-}\sim\{g_{i,j}|o_{i,j}=0\}$.
Figure 4: Overview of iterative guidance optimization. By iteratively updating both the model and the reference policy, Matryoshka progressively refines its intermediate guidance.
3.3 Iterative Guidance Optimization
@@
Supervised Fine-Tuning for Behavior Cloning. To quickly initialize the controller’s policy, we adopt the concept of behavior cloning (BC) from reinforcement learning, which involves learning an initial policy by imitating the actions of an expert agent. This is typically achieved through supervised learning on a set of curated instruction-completion pairs for LLMs. We leverage the capabilities of more advanced models, such as GPT-4 (OpenAI, 2024a), to generate the desired guidance for the black-box LLMs on a small set of samples. This data is then used to perform supervised fine-tuning (SFT) on the white-box LLM controller as an initial warm-up step:
$\mathcal{L}_{\text{SFT}}=-\mathbb{E}_{(x,g)\sim\mathcal{D}_{\text{SFT}}}\left[\sum_{l=1}^{L}\log f_{\theta}(g_{l}|g_{<l},x)\right].$ (4)
@@
Through this SFT process, the white-box LLM controller begins to acquire the capability to effectively guide the subsequent black-box LLM. It can then be utilized to generate high-quality guidance for further optimization steps.
Direct Guidance Optimization. By allowing the warmed-up white-box LLM controller to interact with the black-box LLM environment over multiple turns, we can curate a dataset containing both “good” and “bad” guidance pairs from Matryoshka’s intermediate generations. Following reinforcement learning with human feedback (RLHF) (Bai et al., 2022; Ouyang et al., 2022; Ziegler et al., 2019), we obtain relative feedback by comparing pairs of generated guidance that share the same prompt. The preference signal is modeled using the Bradley-Terry model (Bradley & Terry, 1952). Given an input $x$ and a generated guidance pair $(g^{+},g^{-})$, the model specifies the probability of $g^{+}$ being chosen over $g^{-}$ as:
$p(g^{+}\succ g^{-}|x)=\frac{\exp\left(u\left(g^{+};x\right)\right)}{\exp\left(u\left(g^{+};x\right)\right)+\exp\left(u\left(g^{-};x\right)\right)}=\sigma\left(u\left(g^{+};x\right)-u\left(g^{-};x\right)\right),$ (5)
where $\sigma(x)=\frac{e^{x}}{(e^{x}+1)}$ is the logistic function. This formulation allows us to access sequence-level preferences to optimize the intermediate guidance generated by the white-box LLM controller. Following Rafailov et al. (2024b), we establish a connection between the white-box LLM controller and its associated optimal policy. Specifically, we consider the following KL-regularized planning problem with respect to a reference policy $\pi_{\text{ref}}$:
$\max_{\theta}\mathbb{E}_{x\sim\mathcal{X}}\mathbb{E}_{g\sim\pi_{\theta}(\cdot|x)}\left[u\left(x,g\right)-\eta^{-1}\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\left(g|x\right)||\pi_{\text{ref}}\left(g|x\right)\right]\right].$ (6)
The optimization problem above has a closed-form solution. For any guidance $g$, the optimal policy $\pi^{*}$ is given by:
$\pi^{*}(g|x)\propto\pi_{\text{ref}}(g|x)\exp(\eta u(x,g)).$ (7)
This leads us to the direct preference optimization (DPO) loss for optimizing the intermediate guidance generated by the white-box LLM controller:
$\mathcal{L}_{\text{DPO}}:=\mathbb{E}_{(x,g^{+},g^{-})\sim\mathcal{D}}\left[-\log\sigma\left(\eta^{-1}\left[\log\left(\frac{\pi_{\theta}(g^{+}|x)}{\pi_{\text{ref}}(g^{+}|x)}\right)-\log\left(\frac{\pi_{\theta}(g^{-}|x)}{\pi_{\text{ref}}(g^{-}|x)}\right)\right]\right)\right].$ (8)
@@
Iterative Guidance Optimization. In the previous sections, we discussed sampling positive and negative intermediate generations through multi-turn interactions with the environment. However, an imbalance between positive and negative samples may arise, leading to overfitting on simplistic patterns and hindering the self-improvement of the white-box LLM controller. To address this issue, we propose an iterative guidance optimization method (Figure 4) that interleaves data sampling and training steps. We begin by initializing the model with parameters $\theta^{(0)}=\theta$ without any prior training and sample an initial dataset $\mathcal{D}^{(0)}$ as introduced in Section 3.2. At the $m$-th iteration, we have the optimized model $\theta^{(m)}$. Following STaR (Zelikman et al., 2022), we enhance the model’s generation for the next iteration by bootstrapping the dataset. This involves combining the previous datasets with new samples generated by the current model $\{g_{i,0}^{(m)},g_{i,1}^{(m)},\cdots,g_{i,K\times T}^{(m)}\}\sim f_{\theta^{(m)}}(x)$:
[…]
@@
In the $t$-th iteration, we construct the training dataset $\mathcal{D}^{(t)}$ for DPO by sampling positive and negative pairs for each question. When training the model for the next iteration $\theta^{(m+1)}$, we update the reference policy to be the model from the previous iteration, $\pi_{\theta}^{(m)}(g|x)$. Consequently, the training objective for iterative guidance optimization of the white-box LLM controller becomes:
@@
$\mathcal{L}_{\text{IGO}}:=\mathbb{E}_{(x,g^{+},g^{-})\sim\mathcal{D}}\left[-\log\sigma\left(\eta^{-1}\left[\log\left(\frac{\pi_{\theta}^{(m+1)}(g^{+}|x)}{\pi_{\theta}^{(m)}(g^{+}|x)}\right)-\log\left(\frac{\pi_{\theta}^{(m+1)}(g^{-}|x)}{\pi_{\theta}^{(m)}(g^{-}|x)}\right)\right]\right)\right].$
@@ 4 Experiments
In this section, we present comprehensive experiments on a diverse set of complex, long-horizon tasks to demonstrate the enhanced capabilities of black-box LLMs using Matryoshka.
@@ 4.1 Experimental Setup
Tasks and Datasets. We consider three types of tasks in experiments, each targeting a distinct capability of black-box LLMs: (1) LaMP (Salemi et al., 2023) for personalization capabilities, (2) GSM8K (Cobbe et al., 2021) for reasoning capabilities, and (3) ALFWorld (Shridhar et al., 2020) for planning capabilities. Dataset details are available in Appendix C.
[…]
Evaluation Metrics. For the personalization tasks, consistent with the evaluation metrics specified in LaMP (Salemi et al., 2023), we use accuracy (Acc) and F1 score (F1) for the classification tasks in LaMP-2N and LaMP-2M. For the ordinal multi-class classification task in LaMP-3, we employ mean absolute error (MAE) and root mean squared error (RMSE). To comprehensively evaluate the personalized text generation tasks in LaMP-4 and LaMP-5, we report ROUGE-1 (R-1), ROUGE-L (R-L), and BLEU scores. For the math reasoning task, we assess the models based on the accuracy of obtaining the final correct answer. For the planning task, consistent with previous works (Sun et al., 2024a), we evaluate performance using the success rate (%). The success rate is calculated as the number of successful episodes divided by the total number of episodes. In ALFWorld, an episode is considered a failure if the task remains unsolved after executing 50 actions, which is the maximum allowed number of actions per episode.
Implementations. For the white-box LLM controller, we utilize LLaMA-3-8B-Instruct as the backbone language model. In the black-box LLM environment, our experiments employ gpt-4o-mini for personalization tasks in LaMP, and gpt-3.5-turbo for reasoning and planning tasks in GSM8K and ALFWorld, respectively. All experiments with GPTs are conducted using the Microsoft Azure OpenAI service. Please refer to Appendix E for implementation details.
4.2 Personalization: LaMP
Main Results. Table 1 summarizes the primary experimental results on the LaMP dataset. Our proposed method, Matryoshka, consistently outperforms or matches other state-of-the-art baselines, highlighting its efficacy of advancing black-box LLMs in personalization. For classification tasks, Matryoshka achieves an accuracy of 0.832 on LaMP-2N and 0.535 on LaMP-2M, surpassing other baselines by a significant margin. For generation tasks, Matryoshka also attains over a 25% improvement in BLEU score on LaMP-4. These results demonstrate the effectiveness of Matryoshka in both classification and generative personalization tasks. Furthermore, Matryoshka has the potential to be enhanced with RAG, combining intermediate generations with the retrieved user history data to improve performance further.
@@ 4.2 Personalization: LaMP
gpt-4o-mini (OpenAI, 2024b) 0.514 0.513 0.655 0.473 0.413 0.325 0.371 0.673 0.132 0.116 0.992
RAG (k=1) (Salemi et al., 2023) 0.626 0.624 0.733 0.539 0.444 0.378 0.311 0.631 0.141 0.126 1.296
RAG (k=4) (Salemi et al., 2023) 0.632 0.632 0.792 0.611 0.502 0.430 0.272 0.579 0.161 0.146 2.953
PAG (Richardson et al., 2023) 0.624 0.624 0.775 0.559 0.496 0.443 0.316 0.645 0.143 0.130 1.968
Matryoshka 0.640 0.640 0.832 0.614 0.535 0.475 0.282 0.588 0.171 0.157 4.144
w/o IGO 0.611 0.611 0.807 0.575 0.496 0.432 0.311 0.636 0.131 0.120 1.341
Table 1: Main experimental results on the personalization task using the LaMP benchmark. For all baselines, including Matryoshka, we utilize gpt-4o-mini as the black-box LLM generator. R-1 and R-L refer to ROUGE-1 and ROUGE-L, respectively. $k$ denotes the number of items retrieved. An upward arrow ($\uparrow$) indicates that higher values are preferred, whereas a downward arrow ($\downarrow$) signifies that lower values are better. The best score and second-best score for each task are emphasized in bold and underlined, respectively. IGO represents Intermediate Guidance Optimization. Notations are consistent across tables.
@@
w/o IGO (gemini) 0.568 0.561 0.811 0.602 0.505 0.411 0.365 0.715 0.164 0.150 3.439
Table 2: Plug-and-Play experimental results for gpt-3.5-turbo and gemini-1.5-flash across the LaMP benchmark. We employ Matryoshka pre-trained on gpt-4o-mini as the white-box LLM controller.
Ablation Studies. In our ablation studies, we compare our proposed method, Matryoshka, with a baseline lacking Intermediate Guidance Optimization (IGO) in Table 1. Using the same black-box model (gpt-4o-mini), our optimized white-box controller consistently and significantly outperformed the original LLaMA-3-8B-Instruct. These results demonstrate the effectiveness of IGO in enhancing the white-box controller to generate more informative and higher-quality intermediate outputs, thereby guiding the black-box model toward better final answers.
Plug-and-Play. Matryoshka can seamlessly apply the optimized white-box controller, LLaMA-3-8B-Instruct, to other black-box models in a plug-and-play manner without additional training costs. We further utilize this well-tuned white-box controller as a plug-in to integrate with black-box models such as gpt-3.5-turbo and gemini-1.5-flash. Table 2 presents the plug-and-play results. The experimental results show that our well-tuned controller consistently outperforms other baselines. Specifically, on LaMP-3 and LaMP-4, our plug-in surpasses other baselines by a large margin, demonstrating effectiveness across both classification and generation tasks. The effectiveness of Matryoshka in plug-and-play scenarios arises from the generalization capability of intermediate guidance, which can benefit different black-box LLMs.
(a) LaMP-2M.
(b) LaMP-4.
Figure 5: Effect of number of history per user in LaMP-2M and LaMP-4.
Effects of Profile Count. To further investigate the effect of user profile count on the generation of intermediate outputs, we analyze performance across different numbers of profiles per user. Figure 5 presents the accuracy and ROUGE-L curves separately for LaMP-2M and LaMP-4, with the x-axis representing the total number of profiles per user (e.g., “0-20” indicates users with 0 to 20 profiles). We compared the results of our proposed method, Matryoshka, and PAG, utilizing the white-box controller Llama-3-8B-Instruct and the black-box model gpt-4o-mini. On LaMP-2M, as the profile count increases, PAG’s performance significantly deteriorates, whereas Matryoshka maintains stable performance and surpasses PAG by an increasing margin. For LaMP-4, both Matryoshka and PAG exhibit similar trends, but Matryoshka consistently outperforms PAG by a substantial and steady margin. These results demonstrate the efficacy of IGO in enhancing the summarization capabilities of the black-box controller, especially when dealing with varying and large amounts of profiles.

### v2

@@ 3.2 Data Collection by Interacting with Black-Box LLM Environment
In the MDP formulation, we consider the action space as the set of possible guidance that can enhance the capabilities of black-box LLMs. The observation space is determined by the oracle evaluation function for each task, defined as $\text{eval}(\cdot)$, where the sampled supervision signal is denoted as $z$, with $z=1$ indicating that $f_{\theta}(x)$ is positive guidance while $z=0$ indicating $f_{\theta}(x)$ negative guidance. During the multi-turn interactions, if the observation $o_{t}$ at the $t$-th step returns a negative signal, the next action step $a_{t+1}$ involves modifying the intermediate guidance based on the feedback. The interactions continue until a positive signal is observed or the maximum number of turns $T$ is reached.
For each input $x_{i}$, we perform $T$-step multi-turn interactions with the black-box LLM-based environment to obtain the trajectories $(a_{i,1},o_{i,1},a_{i,2},o_{i,2},\cdots,a_{i,T},o_{i,T})$. To increase the diversity of intermediate generations, we introduce randomness into the policy and repeat the entire interaction process $K$ times. This results in $K$ trajectories, yielding intermediate generations $\{\tau_{i,1},\tau_{i,2},\cdots,\tau_{i,K\times T}\}$ along with their corresponding observations $\{o_{i,1},o_{i,2},\cdots,o_{i,K\times T}\}$, which serve as sampling signals. We then sample the positive trajectory $\tau_{i}^{+}$ from the set of guidance with positive observations, $\tau_{i}^{+}\sim\{\tau_{i,j}|o_{i,j}=1\}$ and the negative trajectory from the remaining generations, $\tau_{i}^{-}\sim\{\tau_{i,j}|o_{i,j}=0\}$.
Figure 4: Overview of iterative guidance optimization. By iteratively updating both the model and the reference policy, M-Pilot progressively refines its intermediate guidance.
3.3 Iterative Direct Preference Optimization
@@
Supervised Fine-Tuning for Behavior Cloning. To quickly initialize the controller’s policy, we adopt the concept of behavior cloning (BC) from reinforcement learning, which involves learning an initial policy by imitating the actions of an expert agent. This is typically achieved through supervised learning on a set of curated instruction-completion pairs for LLMs. We leverage the capabilities of more advanced models, such as GPT-3.5 (schulman2022introducing, ), to generate the desired guidance for the black-box LLMs on a small set of samples. This data is then used to perform supervised fine-tuning (SFT) on the white-box LLM controller as an initial warm-up step:
$\textstyle\mathcal{L}_{\text{SFT}}$ $\textstyle=-\mathbb{E}_{\tau}\!\bigl[\log p(\tau\mid x)\bigr]$ $\textstyle=-\mathbb{E}_{\tau}\!\Bigl[\sum_{t=1}^{T}\log p\bigl(g_{t}\mid x,\,\{g_{\ell}\}_{\ell=1}^{t-1}\bigr)\;\underbrace{\cancel{+\log p\bigl(y\mid x,\,\{g_{t}\}_{t=1}^{T}\bigr)}}_{\text{Non-learnable}}\Bigr].$ (7)
@@
Through this SFT process, the white-box LLM controller begins to acquire the capability to effectively guide the subsequent black-box LLM. It can then be utilized to generate high-quality guidance for further optimization steps. Alternatively, if the initial white-box controller already demonstrates strong performance, we can directly skip the SFT process.
@@
Iterative Preference Pair Collection. By allowing the warmed-up white-box LLM controller to interact with the black-box LLM environment over multiple turns, we can curate a dataset containing both “good” and “bad” guidance pairs from M-Pilot’s intermediate generations. However, an imbalance between positive and negative samples may arise, leading to overfitting on simplistic patterns and hindering the self-improvement of the white-box LLM controller. To address this issue, we propose an iterative guidance optimization method (Figure 4) that interleaves data sampling and training steps. We begin by initializing the model with parameters $\theta^{(0)}=\theta$ without any prior training and sample an initial dataset $\mathcal{D}^{(0)}$ as introduced in Section 3.2. At the $m$-th iteration, we have the optimized model $\theta^{(m)}$. Following STaR (zelikman2022star, ), we enhance the model’s generation for the next iteration by bootstrapping the dataset. This involves combining the previous datasets with new trajectories $\{\tau_{i,0}^{(m)},\tau_{i,1}^{(m)},\cdots,\tau_{i,K\times T}^{(m)}\}$ generated by the current model $\theta^{(m)}$:
$\left\{\begin{aligned} \mathcal{D}_{+}^{(m)}&=\{\tau_{i,j}^{(m)}|o_{i,j}^{(m)}=1\}\cup\mathcal{D}_{+}^{(m-1)},\\ \mathcal{D}_{-}^{(m)}&=\{\tau_{i,j}^{(m)}|o_{i,j}^{(m)}=0\}\cup\mathcal{D}_{-}^{(m-1)}.\end{aligned}\right$. (8)
In the $m$-th iteration, following reinforcement learning with human feedback (RLHF) (bai2022training, ; ouyang2022training, ; ziegler2019fine, ), we construct the training dataset $\mathcal{D}^{(m)}$ by sampling positive and negative generated guidance that share the same prompt.
Iterative DPO. When training the model for the next iteration $\theta^{(m+1)}$, the preference signal is modeled using the Bradley-Terry model (bradley1952rank, ). Given an input $x$ and a generated guidance pair $(g^{+},g^{-})$, the model specifies the probability of $g^{+}$ being chosen over $g^{-}$ as:
$p(\tau^{+}\succ\tau^{-}|x)$ $=\textstyle\frac{\exp\left(u\left(\tau^{+}\right)\right)}{\exp\left(u\left(\tau^{+}\right)\right)+\exp\left(u\left(\tau^{-}\right)\right)}=\sigma\left(u\left(\tau^{+}\right)-u\left(\tau^{-}\right)\right),$ (9)
where $\sigma(x)=\frac{e^{x}}{(e^{x}+1)}$ is the logistic function. This formulation allows us to access sequence-level preferences to optimize the intermediate guidance generated by the white-box LLM controller. Following rafailov2024direct , we establish a connection between the white-box LLM controller and its associated optimal policy. Specifically, we consider the following KL-regularized planning problem with respect to a reference policy $\pi_{\text{ref}}$:
$\max_{\theta}\mathbb{E}_{x}\mathbb{E}_{\tau}\left[u\left(\tau\right)-\eta^{-1}\mathbb{D}_{\text{KL}}\left[\pi_{\theta}\left(\tau|x\right)||\pi_{\text{ref}}\left(\tau|x\right)\right]\right].$
@@
The optimization problem above has a closed-form solution. For any guidance $g$, the optimal policy $\pi^{*}$ is given by $\pi^{*}(g|x)\propto\pi_{\text{ref}}(g|x)\exp(\eta u(x,g))$. To enable the white-box LLM to self-improve, we update the reference policy to be the model from the previous iteration, $\pi_{\text{ref}}=\pi_{\theta}^{(m)}(g|x)$. Consequently, the training objective for iterative guidance optimization of the white-box LLM controller becomes:
@@
$\mathcal{L}_{\text{IDPO}}$ $=\mathbb{E}_{(x,\tau^{+},\tau^{-})\sim\mathcal{D}}\Bigl[-\log\sigma\Bigl(\eta^{-1}\Bigl(\log\Bigl(\tfrac{\pi_{\theta}^{(m+1)}(\tau^{+}|x)}{\pi_{\theta}^{(m)}(\tau^{+}|x)}\Bigr)-\log\Bigl(\tfrac{\pi_{\theta}^{(m+1)}(\tau^{-}|x)}{\pi_{\theta}^{(m)}(\tau^{-}|x)}\Bigr)\Bigr)\Bigr)\Bigr].$ (10)
Taking Eq. 2 into Eq. 10 leads to:
$\tfrac{\pi_{\theta^{(m+1)}}(\tau^{+}|x)}{\pi_{\theta^{(m)}}(\tau^{+}|x)}=\tfrac{p_{\theta^{(m+1)}}\left(\{g_{t}^{+}\}_{t=1}^{T}|x\right)\cancel{p(\hat{y}^{+}|x,\{g_{t}^{+}\}_{t=1}^{T})}}{p_{\theta^{(m)}}\left(\{g_{t}^{+}\}_{t=1}^{T}|x\right)\cancel{p(\hat{y}^{+}|x,\{g_{t}^{+}\}_{t=1}^{T})}}.$ (11)
Similar conclusions can be achieved for $\tfrac{\pi_{\theta^{(m+1)}}(\tau^{-}|x)}{\pi_{\theta^{(m)}}(\tau^{-}|x)}$. Thus, we can rewrite Eq. 10 as:
$\mathcal{L}_{\text{IDPO}}$ $=\mathbb{E}_{(x,\tau^{+},\tau^{-})\sim\mathcal{D}}\Bigl[-\log\sigma\Bigl(\eta^{-1}\Bigl(\log\Bigl(\tfrac{p_{\theta^{(m+1)}}\left(\{g_{t}^{+}\}_{t=1}^{T}|x\right)}{p_{\theta^{(m)}}\left(\{g_{t}^{+}\}_{t=1}^{T}|x\right)}\Bigr)-\log\Bigl(\tfrac{p_{\theta^{(m+1)}}\left(\{g_{t}^{-}\}_{t=1}^{T}|x\right)}{p_{\theta^{(m)}}\left(\{g_{t}^{-}\}_{t=1}^{T}|x\right)}\Bigr)\Bigr)\Bigr)\Bigr].$
@@ 4.1 Experimental Setup
Tasks and Datasets. We consider three types of tasks in experiments, each targeting a distinct capability of black-box LLMs: (1) LaMP (salemi2023lamp, ) for personalization capabilities, (2) GSM8K (cobbe2021training, ) for reasoning capabilities, and (3) ALFWorld (shridhar2020alfworld, ) for planning capabilities. For reasoning, We also test on more challenging MATH hendrycks2021measuring dataset in Appendix B.2, with ablations on outer-loop multi-turn interactions. More tasks details are in Appendix D.
Table 1: Main experimental results on LaMP benchmark. We utilize gpt-4o-mini as the black-box LLM generator for baselines and M-Pilot. R-1 and R-L refer to ROUGE-1 and ROUGE-L, respectively. $k$ denotes the number of items retrieved. $\uparrow$ indicates that higher values are preferred, whereas ($\downarrow$) signifies that lower values are better. The best score and second-best score for each task are emphasized in bold and underlined, respectively. IDPO represents Iterative Direct Preference Optimization. Notations are consistent across tables.
@@
gpt-4o-mini 0.514 0.513 0.655 0.473 0.413 0.325 0.371 0.673 0.132 0.116 0.992
RAG (k=1) (salemi2023lamp, ) 0.626 0.624 0.733 0.539 0.444 0.378 0.311 0.631 0.141 0.126 1.296
RAG (k=4) (salemi2023lamp, ) 0.632 0.632 0.792 0.611 0.502 0.430 0.272 0.579 0.161 0.146 2.953
PAG (richardson2023integrating, ) 0.624 0.624 0.775 0.559 0.496 0.443 0.316 0.645 0.143 0.130 1.968
M-Pilot 0.640 0.639 0.823 0.607 0.527 0.465 0.277 0.581 0.174 0.160 4.298
w/o IDPO 0.611 0.611 0.807 0.575 0.496 0.432 0.311 0.636 0.131 0.120 1.341
[…]
Evaluation Metrics. For personalization tasks, consistent with the evaluation metrics specified in LaMP (salemi2023lamp, ), we use accuracy (Acc) and F1 score (F1) for the classification tasks in LaMP-2N and LaMP-2M. For the ordinal multi-class classification task, LaMP-3, we employ mean absolute error (MAE) and root mean squared error (RMSE). To comprehensively evaluate the personalized text generation tasks in LaMP-4 and LaMP-5, we report ROUGE-1 (R-1), ROUGE-L (R-L), and BLEU scores. For the math reasoning task, we assess the models based on the accuracy of obtaining the final correct answer. For the planning task, consistent with previous works (sun2024adaplanner, ), we evaluate performance using the success rate (%). The success rate is calculated as the number of successful episodes divided by the total number of episodes. In ALFWorld, an episode is considered a failure if the task remains unsolved after executing 50 actions, which is the maximum allowed actions.
Implementations. For the white-box LLM controller, we utilize LLaMA-3-8B-Instruct as the backbone language model, we also consider Qwen2.5-7B-Instruct as the backbone in Appendix B.5. In the black-box LLM environment, our experiments employ gpt-4o-mini for personalization tasks in LaMP, and gpt-3.5-turbo for reasoning and planning tasks in GSM8K and ALFWorld, respectively. Please refer to Appendix F for implementation details.
4.2 Personalization: LaMP
Main Results. Table 1 summarizes the primary experimental results on the LaMP dataset. Our proposed method, M-Pilot, consistently outperforms or matches other state-of-the-art baselines, highlighting its efficacy of advancing black-box LLMs in personalization. For classification tasks, M-Pilot achieves an accuracy of 0.832 on LaMP-2N and 0.535 on LaMP-2M, surpassing other baselines by a significant margin. For generation tasks, M-Pilot also attains over a 25% improvement in BLEU score on LaMP-4. These results demonstrate the effectiveness of M-Pilot in both classification and generative personalization tasks. Furthermore, M-Pilot has the potential to be enhanced with RAG, combining with the retrieved user history data to improve performance.
Table 2: Plug-and-Play results for gpt-3.5-turbo and gemini-1.5-flash across the LaMP benchmark. We employ M-Pilot pre-trained on gpt-4o-mini as the white-box LLM controller.
@@ 4.2 Personalization: LaMP
w/o IDPO (gemini) 0.568 0.561 0.811 0.602 0.505 0.411 0.365 0.715 0.164 0.150 3.439
Plug-and-Play. M-Pilot can seamlessly apply the optimized white-box controller to other black-box models in a plug-and-play manner without additional training costs. We further utilize this well-tuned white-box controller as a plug-in to integrate with black-box models such as gpt-3.5-turbo and gemini-1.5-flash. Table 2 presents the plug-and-play results. The experimental results show that our well-tuned controller consistently outperforms other baselines. Specifically, on LaMP-3 and LaMP-4, our plug-in surpasses other baselines by a large margin, demonstrating effectiveness across both classification and generation tasks. The effectiveness of M-Pilot in plug-and-play scenarios arises from the generalization capability of intermediate guidance, which can benefit different black-box LLMs.
Ablation Studies. For ablation studies on LaMP, we compare our proposed method, M-Pilot, with a baseline lacking Iterative Direct Preference Optimization (IDPO) in Table 1. Using the same black-box model (gpt-4o-mini), our optimized white-box controller consistently and significantly outperformed the original LLaMA-3-8B-Instruct. These results demonstrate the effectiveness of IDPO in enhancing the white-box controller to generate more informative and higher-quality intermediate outputs, thereby guiding the black-box model toward better final answers. Further ablation studies on LaMP are provided in Appendix B.1.